OpenStudy (anonymous):

Solve for y. 15/y+11=10/y+3

OpenStudy (anonymous):

15/y+11=10/y+3 15+11y=10+3y...=11y-3y=10-15 y =8y=-5 =y=-5/8

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

your welcome do you understand how i got the answer?

OpenStudy (anonymous):

yes